OpenStudy (anonymous):

I need help factoring to solve quadratics.2z^2-21z-36?How do I do this?

OpenStudy (anonymous):

\[x = \frac{ -b \pm \sqrt{b^{2}-4ac} }{ 2a }\] Where

OpenStudy (anonymous):

a = 2z b = -21 c= -36

OpenStudy (mathstudent55):

To factor a trinomial of the type ax^2 + bx + c, follow these steps. 1. Multiply ac together. 2. Find two factors of ac that add to b. Call the factors p and q. 3. Break up the middle term, bx, into the two parts, px and qx. 4. Rewrite the trinomial as ax^2 + px + qx + c 5. Factor the ploynomial above by parts. First factor a common factor out of the first two terms. Then factor a common factor out the last two terms. 6. Then factor out the common facotr.

OpenStudy (mathstudent55):

Now that you see the method, let's do it step by step. 1. ac = 2(-36) = -72

OpenStudy (mathstudent55):

2. -72 = (-24)(3); -24 + 3 = -21. Our p & q are -24 and 3

OpenStudy (mathstudent55):

3. -21z = -24z + 3z 4. 2z^2 - 24z + 3z - 36

OpenStudy (mathstudent55):

5. 2z(z - 12) + 3(z - 12) 6. (z - 12)(2z + 3)